OpenStudy (anonymous):

how many chloride ions are present in 3.6 mol of magnesium chloride

OpenStudy (anonymous):

magnesium chloride has formula MgCl2 now just put in relation molar relationship of chlorine and magnesium chloride n(Cl-) = 2 * n(MgCl2) and then use n= N/L -> N=n*L and thats it